The perimeter of a rectangle is to be no greater than 300 in., and the length must be 125 in. Find the maximum width of the rectangle.
Answer by checkley75(3666)

Perimeter=2*length+2*width
300=2*125+2w
300=250+2w
2w=300-250
2w=50
w=50/2
w=25 The maximum measurement of the width.